Word Problem One - An old car has to travels a 2-mile route, uphill and down. Because it is so old, the car can climb the first mile - the ascent - no faster than an average speed of 20 mph. How fast does the car have to travel the second mile - on the descent it can go faster, of course - in order to achieve an average speed of 30 mph for the trip? (Show a complete algebraic solution for the problem and check the answer.)
The teacher said set up a table
Trip DISTANCE = SPEED x TIME
UP
DOWN
TOTAL
I am confused as to when to put in x and how to make an equation!

Answer by rfer(12644)

do an average
30=(20+x)/2
60=20+x
60-20=40 mph
your teacher might be looking for something diferent,
but this works